Vanillin, the dominant flavoring in vanilla, contains C, H,

and O. When 1.05 g of this substance is completely com-
busted, 2.43 g of CO2 and 0.50 g of H2O are produced. What

is the empirical formula of vanillin?
1
Expert's answer
2018-02-15T03:28:44-0500
.m (vanillin)=1.05g
.m (CO2)=2.43 g
.m(H2O)=0.5 g
the empirical formula of vanillin?
Mr (CO2)= Ar (С) +2 Ar (O)= 12+ 2*16=44
Find amount of substances of CO2 : n (CO2)=m(CO2)/Mr(CO2)=2.43/44 =0.055 mol
.n(C) = n(CO2)=0.055 mol
Find amount of substances of H2O: n(H2O)=m(H2O)/Mr(H2O)=0.5/18=0.028
Mr (H2O)= 2Ar(H)+Ar(O)= 2+16=18
.n(H)=2n (H20)= 2*0.028=0.056 mol
.n(C) :n(H)=0.055 :0.056=1:1
the empirical formula is CH